A time study analysis shows an assembly operation for 50 cycles with an average cycle time of 15.00 minutes. If you assigned a performance rating of .95 and a work-day allowance factor of 20 percent, answer the following question:

What is the Observed Time?
Answer #1

Normal Time = (Time worked/No. of units produced) * Performance rating

Normal Time = Average time per unit * Performance rating = 15 * 0.95 = 14.25 minutes

Standard Time = Normal Time / ( 1 - Allowance)

Standard Time = 14.25 / ( 1 - 0.2 ) = 17.8125 Minutes

Observed Time = Standard Time / (Performance rating)(1 + Allowance)

Observed Time = 17.8125 / ( 0.95 * 1.2 ) = 15.625 Minutes